OpenStudy (anonymous):

1. 3y²-3y divided by y-1 2. u+3 divided by u²-9 RESTRICTIONS ON VARIABLES?

OpenStudy (anonymous):

3y²-3y =3y(y-1) and so on... u²-9=u²-3² =(u+9)(u-3)....

OpenStudy (anonymous):

\[(3 y^2-3 y )\div( y-1)=\frac{ 3 y \left( y-1 \right) }{ \left( y-1 \right) }=?\] restrictions \[(y-1)\neq 0~or~y \neq 1\]